1. (3^2) × (2×3^5 )^(-2) × (18)^2;


2. ((2×3/9 ÷3)^(-2)^-1 * (9/4 )^(2 )×(2/5 )^(-1) )


1
Expert's answer
2020-04-20T11:53:52-0400

1. 32 × (2 × 35 )-2 × 182 = 32 × 1/(2 × 35)2 × (2 × 32)2 = (32 × 22 × 34)/(22 × 310) = 1/34 = 3-4 = 1/81


2. (2×3/9 ÷3)-2*(-1) × (9/4 )2 × (2/5)-1 = (2/9)2 × (9/4)2 × 5/2 = (2/9 × 9/4)2 × 5/2 = (1/2)2 × 5/2 = 5/8